The diameter of a sphere is measured to be 4.11in.

(a) Find the radius of the sphere in centimeters.
cm

(b) Find the surface area of the sphere in square centimeters.
cm2

(c) Find the volume of the sphere in cubic centimeters.
cm3
Answer #1

Diameter of sphere = 4.11 in

We know that

1 in = 2.54 cm

So,

Diameter = (4.11 in)*(2.54 cm/1 in)

Diameter = 10.44 cm

Part A.

radius = diameter/2 = 10.44/2 = 5.22 cm

Part B

Surface Area of sphere = 4*pi*r^2

r = radius

SA = 4*pi*5.22^2 = 342.41 cm^2

In 3 sig figures

Surface area = 342 cm^2

Part C.

Volume of sphere = 4*pi*r^3/3

V = 4*pi*5.22^3/3 = 595.80 cm^3

In 3 sig figures

V = 596 cm^3

Please Upvote.